OpenStudy (sleepyjess):

Find the exact solution to the equation. Show your work. \(4^{12-4x}=256\)

OpenStudy (sleepyjess):

@AccessDenied

ganeshie8 (ganeshie8):

hint : 256 = 4^4

OpenStudy (sleepyjess):

so 12-4x must equal 4 right?

ganeshie8 (ganeshie8):

Yup!

ganeshie8 (ganeshie8):

when bases are equal, you can equate the exponents

OpenStudy (sleepyjess):

subtract 12 4x=-8?

OpenStudy (sleepyjess):

divide by 4 x=-2?

ganeshie8 (ganeshie8):

subtract 12 -4x = -8

ganeshie8 (ganeshie8):

divide by 4 -x = -2

ganeshie8 (ganeshie8):

multiply -1 x = 2

OpenStudy (sleepyjess):

thank you so much!